OpenStudy (anonymous):

what is the solution set to the equation x^2+8x+16=49 (((HELP FAST, PLEASE SHOW STEPS!!!!))

OpenStudy (mertsj):

\[(x+4)^2=49\]

OpenStudy (mertsj):

Take square root of both sides.

OpenStudy (anonymous):

Oh, thank you :)

OpenStudy (adilalvi):

Simplifying x2 + 8x + 16 = 49 Reorder the terms: 16 + 8x + x2 = 49 Solving 16 + 8x + x2 = 49 Solving for variable 'x'. Reorder the terms: 16 + -49 + 8x + x2 = 49 + -49 Combine like terms: 16 + -49 = -33 -33 + 8x + x2 = 49 + -49 Combine like terms: 49 + -49 = 0 -33 + 8x + x2 = 0 Factor a trinomial. (-11 + -1x)(3 + -1x) = 0 Subproblem 1 Set the factor '(-11 + -1x)' equal to zero and attempt to solve: Simplifying -11 + -1x = 0 Solving -11 + -1x = 0 Move all terms containing x to the left, all other terms to the right. Add '11' to each side of the equation. -11 + 11 + -1x = 0 + 11 Combine like terms: -11 + 11 = 0 0 + -1x = 0 + 11 -1x = 0 + 11 Combine like terms: 0 + 11 = 11 -1x = 11 Divide each side by '-1'. x = -11 Simplifying x = -11 Subproblem 2 Set the factor '(3 + -1x)' equal to zero and attempt to solve: Simplifying 3 + -1x = 0 Solving 3 + -1x = 0 Move all terms containing x to the left, all other terms to the right. Add '-3' to each side of the equation. 3 + -3 + -1x = 0 + -3 Combine like terms: 3 + -3 = 0 0 + -1x = 0 + -3 -1x = 0 + -3 Combine like terms: 0 + -3 = -3 -1x = -3 Divide each side by '-1'. x = 3 Simplifying x = 3 Solution x = {-11, 3}